# UniTS: A Unified Multi-Task Time Series Model

### v2

@@ 4.2 Unified Network Architecture in UniTS
The complexity of time series samples varies in terms of the number of variables, temporal dynamics, and time lengths across different domains and types of tasks. UniTS uses a modified transformer architecture [101] to handle heterogeneous multi-domain data with varying dynamics and the number of variables (Figure 2c). The following are key modules of UniTS network architecture:
@@
Dylinear. We enhance the transformer block by incorporating a dynamic operator (Dylinear) into the feed-forward network layer (FFN). This modification enables the FFN to capture dependencies between tokens, in contrast to the standard FFN, which processes embedding vectors on a point-wise basis. DyLinear uses a weight interpolation scheme to accommodate varying time lengths. Different variables in a sample share one DyLinear layer. For one variable of a given time series sample of tokens $\mathbf{z}_{t}\in\mathbb{R}^{l_{\text{in}}\times d}$, DyLinear conducts interpolation on the learnable weights $\mathbf{w}\in\mathbb{R}^{w_{\text{in}}\times w_{\text{out}}}$ to accommodate varying time lengths as follows:
$\text{DyLinear}(\mathbf{z}_{t},\mathbf{w})=\mathbf{W_{\text{Interp}}}\mathbf{z}_{t}\ \in\mathbb{R}^{l_{\text{out}}\times d};\mathbf{W_{\text{Interp}}}=\text{Interp}(\mathbf{w})\in\mathbb{R}^{l_{\text{in}}\times l_{\text{out}}},$ (4)
where Interp is a bi-linear interpolation to resize $\mathbf{w}$ from shape $w_{\text{in}}\times w_{\text{out}}$ to $l_{\text{in}}\times l_{\text{out}}$ to match the input and output length. Interpolating weights along the time dimension is intuitive because time series samples, even with different sampling rates, exhibit similar dynamic patterns. DyLinear efficiently captures these relational patterns across various time series samples from diverse datasets, resulting in substantial performance improvements in generative tasks, as demonstrated in Table 19.
Gating module. To mitigate interference in the latent representation space caused by multi-domain and multi-task datasets, we add a gating module after each layer (Figure 2). This module dynamically re-scales features in layer-wise latent spaces and promotes stability of latent representations.
@@
The UniTS architecture includes the backbone network composed of $N$ modified transfomer blocks described above, a CLS tower, and a GEN tower. Implementation details are in Appendix C.3. Ablations in Appendix F verify the effectiveness of this architecture.
@@ 4.3 UniTS Model Training
Unified masked reconstruction pretraining. To enhance UniTS’s abilities to 1) learn general features applicable to both generative and predictive tasks and 2) efficiently adapt to downstream tasks via prompt learning, we introduce a unified mask reconstruction pretraining scheme. It leverages the semantics of both prompt and CLS tokens (Section 4.1) for masked reconstruction pretraining, therefore learning representations for both generative and predictive capabilities. This is distinct from pretraining strategies that use either generative [80, 117, 26, 53] or predictive [71, 106, 114, 29, 121, 85] approach. Unlike these approaches that pretrain only the model backbone, our strategy pretrains all components of UniTS, including the backbone and GEN/CLS towers (Section 4.2), enabling prompt and zero-shot learning over a frozen pre-trained model. Tokens of a full input time series sample $\mathbf{x}$ are randomly masked, i.e., replaced with GEN tokens, and sent to the UniTS backbone network. In the unified pretraining loss, tokens from the backbone network output are sent to the CLS/GEN towers to reconstruct the input sample $\mathbf{x}$, formulating as follows:
$L_{\text{pretrain}}=L_{\text{MSE}}(H_{\texttt{GEN}}(\mathbf{z}_{p},\mathbf{z}_{\mathbf{x}}),\mathbf{x})+L_{\text{MSE}}(H_{\texttt{GEN}}(H_{\texttt{CLS}}(\mathbf{z}_{\text{Pred}}),\ \mathbf{z}_{\mathbf{x}}),\ \mathbf{x}).$ (5)
@@
$L_{\text{MSE}}$ is the MSE loss to predict the full sample $\mathbf{x}$. For the left side of the loss, prompt token $\mathbf{z}_{p}$ is sent along with sample token $\mathbf{z}_{\mathbf{x}}$ to GEN tower $H_{\texttt{GEN}}$ to help with the reconstruction. For the right side of the loss, to leverage the semantics of the CLS token and train the CLS tower $H_{\texttt{CLS}}$ for predictive tasks, $\mathbf{z}_{\text{Pred}}$ (Eq. 2) from the model output is processed by the CLS tower $H_{\texttt{CLS}}$ to get classification-related embedding vectors $\mathbf{\hat{z}}_{\text{Pred}}=H_{\texttt{CLS}}(\mathbf{z}_{\text{Pred}})$, and an GEN tower $H_{\texttt{GEN}}$ takes in $\mathbf{\hat{z}}_{\text{Pred}}$ and $\mathbf{z}_{\mathbf{x}}$ to predict the full sample. This unified pretraining strategy involves pretraining both tokens, the backbone network, and the GEN/CLS towers for both generative and predictive abilities.
@@
Training UniTS models. We implement and evaluate two UniTS models, each trained in a different regime. We start with a pre-trained UniTS that is optimized using self-supervised $L_{\text{pretrain}}$ in Eq. 5 and trained across a collection of multi-domain datasets. Given a self-supervised pre-trained UniTS whose weights are frozen, we consider a fine-tuned model where only tokens for predictive or generative tasks are fine-tuned (denoted as UniTS-PMT in Experiments). We also consider a standard multi-task supervised learning regime, where a single UniTS model is trained from scratch to simultaneously perform many tasks (denoted as UniTS-SUP in Experiments). These two regimes all use a multi-task setup, where a single model is trained and tested on multiple tasks and datasets. During multi-task training, we sample batches of time series samples and aggregate dataset-centric loss values: $L_{\text{total}}=\sum_{i=1}^{I}\lambda_{i}L_{i}(D_{i})$, where $L_{i}$ is the loss of batch $i$, $\lambda_{i}$ is the weight for each loss, and $I$ denotes the number of batches. We follow [109] and use the MSE loss for forecasting and cross-entropy loss for classification.
Additionally, for fair comparison with models trained in a single-task manner, we follow the experimental setup of [109, 66] and benchmark UniTS in a single-task setting (denoted as UniTS-ST in Experiments), where every model is trained and evaluated separately on each dataset/task.
@@ 5.1 Benchmarking UniTS on Single-Task Learning
Setup. To make fair comparisons with time series baselines, we benchmark the single-task performance of UniTS-ST on forecasting, classification, anomaly detection, and imputation. Models are separately trained from scratch with different configurations per dataset. Details are in Appendix K.
Results. Table 1 shows the single-task performance for four types of tasks. On forecasting tasks with forecasting lengths of 92, 196, 336, and 720, compared with 11 forecasting methods, UniTS-ST achieves the best results on 32 out of 36 datasets for MSE and 27 out of 36 for MAE, surpassing the previous best method, iTransformer, by a clear margin. In Table 29, we demonstrate that UniTS-ST outperforms the concurrent MOMENT [36] model, which was trained on a large and diverse collection of time series data. Additionally, UniTS-ST achieves stronger performance than LLM-reprogrammed methods that are pre-trained with extensive natural language data, e.g. GPT4TS [126], TEST [95], LLM4TS [12], and TEMPO [10]. On 10 classification datasets, UniTS-ST outperforms 19 classification methods on the average accuracy, such as the transformer/MLP/frequency-based methods. It has the gain of 1.4% compared to the previous best TimesNet model. On 5 anomaly detection datasets, UniTS-ST has a clear gain of 3.95% in F1 score compared to the TimesNet, and also beat other 15 anomaly detection methods, such as Anomaly Transformer [113]. On 16 imputation datasets with mask ratio of 12.5%, 25%, 37.5%, UniTS-ST has the best results on all datasets in terms of MSE and MAE, outperforming 14 baseline methods. UniTS-ST has the SoTA performance on these single-task benchmarks, showing its effectiveness.
@@
Table 2: Multi-task benchmarking across 20 forecasting tasks and 18 classification tasks. Both UniTS-SUP and UniTS-PMT process all 38 tasks using a single model. GPT4TS reprograms a pre-trained LLM (GPT-2) to time series and has dataset/task-specific modules. “P” is forecasting length. “Class./Num.” denotes the “number of classes in each task”/“number of datasets”.
@@ 5.2 Benchmarking UniTS for Multi-Task Learning
Setup. In a multi-task setting, we benchmark a single UniTS model co-trained and evaluated on 38 datasets, comprising 20 forecasting tasks and 18 classification tasks, with variations in the number of variables/sensors, classification classes, and forecasting lengths. We consider two variants of UniTS; the fully supervised UniTS-SUP and the more challenging UniTS-PMT with prompting, as introduced in Section 4.3. Baselines use the same fully supervised multi-task training as our approach, but cannot handle differences across data types and task specifications with a single model. To benchmark them, a shared backbone is used for all tasks, augmented by data-specific input modules and task-specific output modules.
Figure 3: Zero-shot forecasting to new forecasting lengths. UniTS achieves any new forecasting length with unified direct multi-step inference. Baseline methods use the sliding windows inference as they do not support direct multi-step inference.
@@
Results: benchmarking of UniTS. Table 2 shows multi-task learning performance. UniTS consistently outperforms baseline methods, achieving the best results in 17 out of 20 forecasting tasks (MSE) and 10 out of 18 classification tasks (accuracy). Performance gains are especially remarkable because UniTS has one fully shared model, whereas all existing methods require task or dataset-specific modules. We find that baseline methods encounter difficulties performing well across different types of tasks. For example, TimesNet, which excels in classification tasks, underperforms in forecasting tasks. Conversely, iTransformer, the top-performing forecaster, struggles with classification tasks. In contrast, the UniTS model exhibits robust performance across classification and forecasting. On forecasting, UniTS-SUP surpasses the leading baseline, iTransformer, by 5.8% (0.439 vs. 0.466) in MSE and 3.3% (0.381 vs. 0.394) in MAE. On classification, UniTS-SUP has an average gain of 0.7% accuracy (81.6% vs. 80.9%) over the strongest baseline (TimesNet). UniTS shows promising potential to unify data and task diversity across time series domains.
@@
Results: prompt learning is competitive with supervised training. Using tokens to prompt a frozen UniTS, SSL-pretrained UniTS achieves performance comparable to supervised trained UniTS (Table 2). Notably, UniTS-PMT surpasses supervised learning on forecasting, as indicated by a lower MAE score (0.379 vs. 0.381), suggesting effectiveness of prompt learning in UniTS. Additionally, prompt learning with UniTS already exceeds the performance of supervised baseline methods with separate modules. This indicates that the SSL-pretrained model learns valuable representations for time series tasks, and prompt learning allows the model to efficiently adapt to target tasks.
5.3 UniTS for Direct Multi-Step Forecasting to Varying Time Lengths
Setup. Direct multi-step forecasting is tasked with predicting new lengths by adjusting from the trained length, with offsets ranging from 0 to 384 in our setting. We keep 14 out of 20 forecasting datasets with a wide range of lengths. UniTS can easily achieve this task by simply repeating the GEN token as described in Section 4.1, which is not supported by existing methods. For comparison with baselines, we have developed a sliding-window forecasting scheme that predicts over a fixed window size and then slides to accommodate new lengths.
Results: direct multi-step inference in UniTS outperforms sliding-window inference. In Figure 3, UniTS demonstrates a notable improvement in performance compared to baselines across various forecasting lengths when using the sliding-window approach. For instance, in the case of the longest forecasting extension of +384, UniTS surpasses the iTransformer by 8.7% in MSE, scoring 0.451 against 0.494. When applying the direct multi-step inference, UniTS achieves an even greater advantage over the iTransformer, improving MSE by 10.5% (0.442 vs. 0.494). This approach significantly reduces the average number of inference steps from 3.66 to 1, resulting in a roughly 3× acceleration in inference steps.
@@ 5.4 UniTS for Few-Shot Learning on New Datasets and Tasks
For the transfer learning on new tasks and datasets, we load the model weights pre-trained on 38 datasets and apply the multi-task setting. We conduct either fully fine-tuning (UniTS-FT) to fine-tune all weights or prompt learning (UniTS-PMT) to train only tokens for new tasks and datasets.
Few-shot classification and forecasting: setup. Pre-trained models, undergo fine-tuning using 5%, 15%, and 20% of the 11 training set shown in Table 8. Average performance are reported as metrics.
@@
Few-shot imputation: setup. Models are fine-tuned with 10% of 6 imputation training data listed in Table 10, asked to impute 25% and 50% of missing data points.
Results. As depicted in Table 4, UniTS-FT with one model significantly surpasses baseline models that employ separate task-specific modules. This indicates UniTS has robust few-shot learning capability in the imputation task. Specifically, on a 25% masking ratio, UniTS-FT exceeds the top-performing baseline iTransformer by 12.4% in MSE and 7.9% in MAE. The margin remains notable at a 50% masking ratio, where UniTS-FT surpasses iTransformer by 8.8% in MSE and 6.8% in MAE. Notably, UniTS-PMT, the fixed model with appropriate prompt tokens not only outperforms all baseline methods but also achieves comparable results to its fully fine-tuned counterpart. This demonstrates that prompting alone can effectively adapt UniTS for imputation.
Few-shot anomaly detection: setup. The pre-trained models have been fine-tuned using 5% of five training dataset as listed in Table 10. Average F1-score is used as the metric.
Results. UniTS outperforms the top-performing baseline (PathTST) across all metrics (Table 5). UniTS-FT achieves an F1-score of 86.3 compared to PathTST’s F1-score of 84.3. UniTS-PMT also outperforms specialized baselines (Anomaly Transformer) trained from scratch.
Additional results and ablations. We explore the zero-shot forecasting ability of UniTS on new datasets in Appendix G. Additional analysis and ablation results are in Appendix F and Appendix E.

### v3

@@ 4.2 Unified Network Architecture in UniTS
Time series samples can have varying numbers of variables, temporal dynamics, and time lengths across different domains and types of tasks. UniTS uses a modified transformer architecture [104] to handle heterogeneous multi-domain data with varying dynamics and the number of variables (Figure 2c). In the following, we describe key modules of UniTS architecture. Note that UniTS can also be used with other backbones, such as Mamba [38].
@@
DyLinear. We modify the transformer block by adding a dynamic operator (DyLinear) into the feed-forward network layer (FFN). This modification enables the FFN to capture dependencies between tokens. In contrast to the standard FFN, which processes embedding vectors on a point-wise basis, DyLinear uses weight interpolation to accommodate varying time lengths. Given a sequence of sample tokens $\mathbf{z}_{t}\in\mathbb{R}^{l_{\text{in}}\times d}$, DyLinear interpolates weights $\mathbf{w}\in\mathbb{R}^{w_{\text{out}}\times w_{\text{in}}}$ to accommodate varying time lengths as follows:
$\text{DyLinear}(\mathbf{z}_{t},\mathbf{w})=\mathbf{W_{\text{Interp}}}\mathbf{z}_{t}\ \in\mathbb{R}^{l_{\text{out}}\times d};\mathbf{W_{\text{Interp}}}=\text{Interp}(\mathbf{w})\in\mathbb{R}^{l_{\text{out}}\times l_{\text{in}}},$ (4)
where Interp is a bi-linear interpolation to resize $\mathbf{w}$ from shape $w_{\text{out}}\times w_{\text{in}}$ to $l_{\text{out}}\times l_{\text{in}}$ to match the input and output length. DyLinear captures dependency patterns across time series samples, which leads to improved performance on generative tasks (Table 23).
Gating module. We add a gating module after each layer to mitigate interference in the latent representation space caused by multi-domain and multi-task datasets (Figure 2). This module dynamically re-scales features in layer-wise latent spaces and promotes the stability of latent representations.
@@
The UniTS architecture includes the backbone network composed of $N$ modified transformer blocks described above, a CLS tower, and a GEN tower. Implementation details are in Appendix C.3. Ablations in Appendix F verify the effectiveness of this architecture.
@@ 4.3 UniTS Model Training
Unified masked reconstruction pre-training. To enhance UniTS’s abilities to 1) learn general features applicable to both generative and predictive tasks and 2) efficiently adapt to downstream tasks via prompt learning, we introduce a unified mask reconstruction pre-training scheme. It leverages the semantics of both prompt and CLS tokens (Section 4.1) for masked reconstruction pre-training, therefore learning representations for both generative and predictive capabilities. This is distinct from pre-training strategies that use either generative [82, 120, 26, 54] or predictive [72, 109, 117, 29, 124, 87] approach. Unlike these approaches that pre-train only the model backbone, our strategy pre-trains all components of UniTS, including the backbone and GEN/CLS towers (Section 4.2), enabling prompt and zero-shot learning over a frozen pre-trained model. For each time-series sample $\mathbf{x}$, a handful of sample tokens get masked and replaced with GEN tokens. These masked sample tokens is then concatenated with prompt tokens and CLS tokens, sent to the UniTS backbone network. In the unified pre-training loss, tokens from the backbone network output are sent to the CLS/GEN towers to reconstruct the input sample $\mathbf{x}$, formulating as follows:
$L_{\text{pretrain}}=L_{\text{MSE}}(H_{\texttt{GEN}}(\mathbf{z}_{p},\mathbf{z}_{\mathbf{x}}),\mathbf{x})+L_{\text{MSE}}(\hat{H}_{\texttt{GEN}}(H_{\texttt{CLS}}(\mathbf{z}_{\text{Pred}}),\ \mathbf{z}_{\mathbf{x}}),\ \mathbf{x}).$ (5)
@@
$L_{\text{MSE}}$ is the MSE loss to predict the full sample $\mathbf{x}$. For the left side of the loss, prompt token $\mathbf{z}_{p}$ is sent along with sample token $\mathbf{z}_{\mathbf{x}}$ to GEN tower $H_{\texttt{GEN}}$ to help with the reconstruction. For the right side of the loss, to leverage the semantics of the CLS token and train the CLS tower $H_{\texttt{CLS}}$ for predictive tasks, $\mathbf{z}_{\text{Pred}}$ (Eq. 2) from the model output is processed by the CLS tower $H_{\texttt{CLS}}$ to get classification-related embedding vectors $\mathbf{\hat{z}}_{\text{Pred}}=H_{\texttt{CLS}}(\mathbf{z}_{\text{Pred}})$, and another GEN tower $\hat{H}_{\texttt{GEN}}$ takes in $\mathbf{\hat{z}}_{\text{Pred}}$ and $\mathbf{z}_{\mathbf{x}}$ to predict the full sample. $\hat{H}_{\texttt{GEN}}$ is only used for pre-training and will be removed for downstream tasks. This unified pre-training strategy involves pre-training both tokens, the backbone network, and the GEN/CLS towers for both generative and predictive abilities.
@@
Training UniTS models. We implement and evaluate two UniTS models, each trained in a different regime. We start with a pre-trained UniTS that is optimized using self-supervised $L_{\text{pretrain}}$ in Eq. 5 and trained across a collection of multi-domain datasets. Given a self-supervised pre-trained UniTS whose weights are frozen, we consider a fine-tuned model where only tokens for predictive or generative tasks are fine-tuned (denoted as UniTS-PMT in Experiments). We also consider a standard multi-task supervised learning regime, where a single UniTS model is trained from scratch to simultaneously perform many tasks (denoted as UniTS-SUP in Experiments). These two regimes use a multi-task setup, where a single model is trained and tested on multiple tasks and datasets. During multi-task training, we sample batches of time series samples and aggregate dataset-centric loss values: $L_{\text{total}}=\sum_{i=1}^{I}\lambda_{i}L_{i}(D_{i})$, where $L_{i}$ is the loss of batch $i$, $\lambda_{i}$ is the weight for each loss, and $I$ denotes the number of batches. We follow [112] and use the MSE loss for forecasting and cross-entropy loss for classification. For fair comparison with models trained in a single-task manner, we follow the experimental setup of [112, 67] and benchmark UniTS in a single-task setting (denoted as UniTS-ST in Experiments), where the model is trained separately on each dataset/task.
@@ 5.1 Benchmarking UniTS on Single-Task Learning
Setup. For fair comparisons with baseline methods, we benchmark single-task UniTS-ST on forecasting, classification, anomaly detection, and imputation. Models are separately trained from scratch with configuration tailored to datasets. Details are in Appendix K.
Results. Table 1 shows the single-task performance for four types of tasks. On forecasting tasks with forecasting lengths of 92, 196, 336, and 720, compared with 11 forecasting methods, UniTS-ST achieves the best results on 28 out of 32 datasets for MSE and 27 out of 32 for MAE, surpassing the previous best method, iTransformer, by a clear margin. In Table 34, we demonstrate that UniTS-ST outperforms the concurrent MOMENT [36] model, which was trained on a large and diverse collection of time series data. Additionally, UniTS-ST achieves stronger performance than LLM-reprogrammed methods that are pre-trained with extensive natural language data, e.g. GPT4TS [129], TEST [97], LLM4TS [12], and TEMPO [10]. On 10 classification datasets, UniTS-ST outperforms 19 classification methods on the average accuracy, such as the transformer/MLP/frequency-based methods. It has a gain of 1.4% compared to the previous best TimesNet model. On 5 anomaly detection datasets, UniTS-ST has a clear gain of 3.95% in F1 score compared to the TimesNet and also beat other 15 anomaly detection methods, such as Anomaly Transformer [116]. On 16 imputation datasets with a mask ratio of 12.5%, 25%, 37.5%, UniTS-ST has the best results on all datasets in terms of MSE and MAE, outperforming 14 baseline methods. UniTS-ST has the SoTA performance on these single-task benchmarks, showing its effectiveness.
@@
Table 2: Multi-task benchmarking across 20 forecasting tasks and 18 classification tasks. Both UniTS-SUP and UniTS-PMT process all 38 tasks using a single model. GPT4TS reprograms a pre-trained LLM (GPT-2) to time series and has dataset/task-specific modules, thus, it is excluded from best count evaluations to ensure fair comparisons.
P” is forecasting length. “Class./Num.” denotes the “number of classes in each task”/“number of datasets”.
@@ 5.2 Benchmarking UniTS for Multi-Task Learning
Setup. In a multi-task setting, we benchmark a single UniTS model co-trained and evaluated on 38 datasets, comprising 20 forecasting tasks and 18 classification tasks, with variations in the number of variables/sensors, classification classes, and forecasting lengths. We consider two variants of UniTS; the fully supervised UniTS-SUP and the more challenging UniTS-PMT with prompting, as introduced in Section 4.3. Baselines use the same fully supervised multi-task training as our approach but cannot handle differences across data types and task specifications with a single model. To benchmark them, a shared backbone is used for all tasks, augmented by data-specific input modules and task-specific output modules.
Figure 3: Direct multi-step forecasting on new lengths. UniTS achieves any new forecasting length with unified direct multi-step inference. Baseline methods use the sliding windows inference as they do not support direct multi-step inference.
@@
Results: Model benchmarking. Table 2 shows multi-task learning performance. UniTS consistently outperforms baseline methods, achieving the best results in 17 out of 20 forecasting tasks (MSE) and 10 out of 18 classification tasks (accuracy). Performance gains are especially remarkable because UniTS has one fully shared model, whereas all existing methods require task or dataset-specific modules. We find that baseline methods encounter difficulties performing well across different types of tasks. For example, TimesNet, which excels in classification tasks, underperforms in forecasting tasks. Conversely, iTransformer, the top-performing forecaster, struggles with classification tasks. In contrast, the UniTS model exhibits robust performance across classification and forecasting. On forecasting, UniTS-SUP surpasses the leading baseline, iTransformer, by 5.8% (0.439 vs. 0.466) in MSE and 3.3% (0.381 vs. 0.394) in MAE. On classification, UniTS-SUP has an average gain of 0.7% accuracy (81.6% vs. 80.9%) over the strongest baseline (TimesNet). UniTS shows promising potential to unify data and task diversity across time series domains.
@@
Results: Prompting is competitive with supervised training. Using tokens to prompt a frozen UniTS, the SSL-pre-trained UniTS achieves performance comparable to its fully supervised counterpart (Table 2). UniTS-PMT even outperforms the supervised model in forecasting, with a lower MAE score (0.379 vs. 0.381), highlighting the effectiveness of prompt learning in UniTS. Furthermore, prompt learning with UniTS surpasses the performance of supervised baseline methods with separate modules. This indicates that the SSL-pre-trained model captures valuable time series representations and that prompt learning allows the model to efficiently adapt to target tasks.
5.3 UniTS for Direct Multi-Step Forecasting
Setup. Direct multi-step forecasting predicts across varying time horizons by adjusting from the original trained length, with offsets ranging from 0 to 384. We use 14 out of 20 forecasting datasets with varying lengths. UniTS achieves this flexibility by repeating the GEN token, as described in Section 4.1, a capability not supported by existing methods. For comparison with baseline models, we implement a sliding-window approach for forecasting. In this method, predictions are made over a fixed window size, which then shifts forward incrementally to cover progressively extended time horizons. This sliding mechanism allows us to adapt the model to forecast over new, unseen time periods while maintaining consistency with the evaluation setup used by baseline methods.
Results: Direct multi-step inference outperforms sliding window approach. In Figure 3, UniTS demonstrates improved performance over baseline models across various forecasting lengths when using the sliding-window approach. For example, in the longest forecasting extension of +384, UniTS outperforms the iTransformer by 8.7% in MSE, achieving a score of 0.451 compared to 0.494. When using direct multi-step inference, UniTS gains an even larger advantage over the iTransformer, reducing MSE by 10.5% (0.442 vs. 0.494). This approach also reduces the average number of inference steps from 3.66 to 1, resulting in a 3× speedup.
@@ 5.4 UniTS for Few-Shot Learning on New Datasets and Tasks
For transfer learning on new tasks and datasets, we load the model weights pre-trained on 38 datasets and apply them in a multi-task setting. We evaluate two approaches: the fully fine-tuned UniTS-FT model and the prompted UniTS-PMT model, in which task-specific tokens are trained.
Setup: Few-shot classification and forecasting. Pre-trained models, undergo fine-tuning using 5%, 15%, and 20% of the 11 training set shown in Table 8. Average performance is reported.
@@
Setup: Few-shot imputation. Models are fine-tuned with 10% of 6 imputation training data listed in Table 10, asked to impute 25% and 50% of missing data points.
Results. A unified UniTS-FT outperforms models that use separate task-specific modules (Table 4), indicating that UniTS has robust few-shot imputation performance. Specifically, on a 25% masking ratio, UniTS-FT exceeds the top-performing baseline iTransformer by 12.4% in MSE and 7.9% in MAE. The margin remains notable at a 50% masking ratio, where UniTS-FT surpasses iTransformer by 8.8% in MSE and 6.8% in MAE. UniTS-PMT, the fixed model with appropriate prompt tokens, outperforms all baseline methods and achieves results comparable to its fully fine-tuned counterpart, suggesting that prompting can adapt UniTS for imputation.
Setup: Few-shot anomaly detection. The pre-trained models have been fine-tuned using 5% of five training datasets as listed in Table 10. The average F1-score is used as the metric.
Results. UniTS outperforms the top-performing baseline (PathTST) across all metrics (Table 5). UniTS-FT achieves an F1-score of 86.3 compared to PathTST’s F1-score of 84.3. UniTS-PMT also outperforms specialized models (Anomaly Transformer) trained from scratch.
Additional results and ablations. Zero-shot learning is significantly more challenging than few-shot learning. Our work primarily focuses on few-shot learning, with some initial exploration of zero-shot learning for forecasting tasks of UniTS on new datasets in Appendix G. Additional analysis and ablation results are in Appendix F and Appendix E.